# LRANet++: Low-Rank Approximation Network for Accurate and Efficient Text Spotting

### v1

@@ Abstract
End-to-end text spotting aims to jointly optimize text detection and recognition within a unified framework. Despite significant progress, designing an accurate and efficient end-to-end text spotter for arbitrary-shaped text remains largely unsolved. We identify the primary bottleneck as the lack of a reliable and efficient text detection method. To address this, we propose a novel parameterized text shape method based on low-rank approximation for precise detection and a triple assignment detection head to enable fast inference. Specifically, unlike other shape representation methods that employ data-irrelevant parameterization, our data-driven approach derives a low-rank subspace directly from labeled text boundaries. To ensure this process is robust against the inherent annotation noise in this data, we utilize a specialized recovery method based on an $\ell_{1}$-norm formulation, which accurately reconstructs the text shape with only a few key orthogonal vectors. By exploiting the inherent shape correlation among different text contours, our method achieves consistency and compactness in shape representation. Next, the triple assignment scheme introduces a novel architecture where a deep sparse branch (for stabilized training) is used to guide the learning of an ultra-lightweight sparse branch (for accelerated inference), while a dense branch provides rich parallel supervision. Building upon these advancements, we integrate the enhanced detection module with a lightweight recognition branch to form an end-to-end text spotting framework, termed LRANet++, capable of accurately and efficiently spotting arbitrary-shaped text. Extensive experiments on several challenging benchmarks demonstrate the superiority of LRANet++ compared to state-of-the-art methods. Code will be available at: https://github.com/ychensu/LRANet-PP.git
@@ 4.1 Datasets
Synth150K [liu2020abcnet] is a synthetic dataset consisting of $54,327$ curved text images and $94,723$ multi-oriented images. Total-Text [ch2017total] is an arbitrary-shaped word-level scene text dataset. It contains $1,255$ images, with $1,000$ for training and $255$ for testing. Each image contains at least one example of curved text. CTW1500 [liu2019curved] is another important arbitrary-shaped scene text benchmark, containing $1,000$ training and $500$ test images. MSRA-TD500 [yao2012detecting] is a multi-language text detection dataset that consists of $300$ training images and $200$ test images, with annotations in a text-line level format. ICDAR 2013 (IC13) [karatzas2013icdar] is a horizontal dataset that contains $229$ training and $233$ test images. ICDAR 2015 (IC15) [karatzas2015icdar] is a multi-oriented scene text dataset that includes $1,000$ training and $500$ test images. The text instances are labeled at the word-level. ICDAR17 MLT (MLT17) [nayef2017icdar2017] is a multi-language scene text dataset containing $9,000$ training images. Following [ye2023deepsolo, huang2023estextspotter], we use the English samples from the training set to pre-train our text spotting model. TextOCR [singh2021textocr] is currently the largest real dataset for text spotting. It is composed of $21,749$ training, $3,153$ validation, and $3,232$ test images. Inverse-Text [ye2023dptext] is a recently proposed dataset focused on inverse-like scenes. It consists of $500$ images for testing, with about 40% being inverse-like instances. ICDAR19 ArT [chng2019icdar2019] is currently the largest arbitrary-shaped dataset, formed by combining and extending CTW1500 and Total-Text. It contains $5,603$ training images. ICDAR19 LSVT [sun2019icdar] is a large-scale Chinese scene text dataset with 30,000 training images. ReCTS [zhang2019icdar] is a benchmark for Chinese text on signboards, containing 20,000 training and 5,000 test images. VinText [nguyen2021dictionary] is a Vietnamese text dataset, including 1,200 training and 500 testing images.
@@ 4.2 Implemented Details
In the testing stage, we set the short sides of the test images to $736$, $1000$, $1000$, $800$, $1440$, $1056$, $1056$ for CTW1500, Total-Text, Inverse-Text, MSRA-TD500, and ICDAR 2015, ReCTS, and VinText, respectively. Following [liu2021abcnet, ye2023dptext, wang2022tpsnet], we adopt IoU@0.5 as the evaluation metric for text detection across all datasets. The evaluation metric for text spotting is consistent with [liu2020abcnet, huang2022swintextspotter, huang2023estextspotter]. The model is trained distributively on 4 NVIDIA RTX 3090 GPUs with a batch size of 4 per GPU, and all the inference speeds listed in the tables are tested on a single NVIDIA RTX 3090 GPU. When using the $250k$ iteration pre-training schedule for text spotting, the whole training process takes about 2.5 days. Recall, Precision, and F-measure are omitted as R”, P”, and “F”, respectively.
@@ 4.3.1 Different Data Representations
TABLE III: Robustness comparison of FMS and SVD on CTW1500, where the CTW1500 training labels are corrupted with 20% artificial noise.
Method Detection E2E IoU
@@
SVD 84.5 87.5 86.0 66.7 $93.6^{(\downarrow 4.4)}$
@@ 4.3.3 Different Low-Rank Subspace Projections
We adopt FMS instead of SVD to estimate the underlying low-dimensional subspace of labeled text contours, as FMS is theoretically more robust to the outliers often found in real-world annotations. We conduct two experiments to verify the advantages of this choice.
First, to validate the necessity of a robust method against the annotation errors inevitably introduced in real-world datasets, we create a corrupted version of the CTW1500 training set by injecting 20% spike noise, where one to five random vertices are drastically displaced. We then learn the LRA basis from this noisy data. As shown in Table III, SVD’s representation quality collapses under noise, dropping 4.4%. In contrast, our FMS method remains stable, dropping by only 0.9%. This robust performance gives FMS a commanding 3.6% lead in IoU on the noisy dataset. Consequently, FMS outperforms SVD by 0.8% and 0.6% in the F-measure for text detection and spotting, respectively.
Second, we compare their performance on geometrically complex shapes using the highly curved text subset from [zhu2021fourier]. As shown in Table IV, FMS builds a higher quality projection. It gives an IoU gain of 2.5% compared to SVD when the projection dimension is low (e.g., 6). Consequently, the FMS method achieves an increase of 0.7% and 0.8% in F-measure for text detection and spotting, respectively. Even with only 6 dimensions for text shape representation, the FMS method still demonstrates good spotting performance on CTW1500 (62.4% in terms of F-measure). When the dimension further increases to 14, a prominent IoU improvement of 10.1% is observed, still better than the SVD method by 0.3%. This suggests our LRA method has promising potential as a compact and effective shape representation module for various tasks.

### v2

@@ Abstract
End-to-end text spotting aims to jointly optimize text detection and recognition within a unified framework. Despite significant progress, designing an accurate and efficient end-to-end text spotter for arbitrary-shaped text remains challenging. We identify the primary bottleneck as the lack of a reliable and efficient text detection method. To address this, we propose a novel parameterized text shape representation based on low-rank approximation for precise detection and a triple assignment detection head for fast inference. Specifically, unlike current data-irrelevant shape representation methods, we exploit shape correlations among labeled text boundaries to construct a robust low-rank subspace. By minimizing an $\ell_{1}$-norm objective, we extract orthogonal vectors that capture the intrinsic text shape from noisy annotations, enabling precise reconstruction via the linear combination of only a few basis vectors. Next, the triple assignment scheme decouples training complexity from inference speed. It utilizes a deep sparse branch to guide an ultra-lightweight inference branch, while a dense branch provides rich parallel supervision. Building upon these advancements, we integrate the enhanced detection module with a lightweight recognition branch to form an end-to-end text spotting framework, termed LRANet++, capable of accurately and efficiently spotting arbitrary-shaped text. Extensive experiments on challenging benchmarks demonstrate the superiority of LRANet++ compared to state-of-the-art methods. Code is available at: https://github.com/ychensu/LRANet-PP.
@@ 4.1 Datasets
Synth150K [liu2020abcnet] is a synthetic dataset consisting of $54,327$ curved text images and $94,723$ multi-oriented images. Total-Text [ch2017total] is an arbitrary-shaped word-level scene text dataset. It contains $1,255$ images, with $1,000$ for training and $255$ for testing. Each image contains at least one example of curved text. CTW1500 [liu2019curved] is another important arbitrary-shaped scene text benchmark, containing $1,000$ training and $500$ test images. MSRA-TD500 [yao2012detecting] is a multi-language text detection dataset that consists of $300$ training images and $200$ test images. ICDAR 2013 (IC13) [karatzas2013icdar] is a horizontal dataset that contains $229$ training and $233$ test images. ICDAR 2015 (IC15) [karatzas2015icdar] is a multi-oriented scene text dataset that includes $1,000$ training and $500$ test images. The text instances are labeled at the word-level. ICDAR17 MLT (MLT17) [nayef2017icdar2017] is a multi-language scene text dataset containing $9,000$ training images. TextOCR [singh2021textocr] is currently the largest real dataset for text spotting. It is composed of $21,749$ training, $3,153$ validation, and $3,232$ test images. Inverse-Text [ye2023dptext] is a recently proposed dataset focused on inverse-like scenes. It consists of $500$ images for testing, with about 40% being inverse-like instances. ICDAR19 ArT [chng2019icdar2019] is currently the largest arbitrary-shaped dataset. It contains $5,603$ training images. ICDAR19 LSVT [sun2019icdar] is a large-scale Chinese scene text dataset with 30,000 training images. ReCTS [zhang2019icdar] is a benchmark for Chinese text on signboards, containing 20,000 training and 5,000 test images. VinText [nguyen2021dictionary] is a Vietnamese text dataset, including 1,200 training and 500 testing images.
@@ 4.2 Implemented Details
In the testing stage, we set the short sides of the test images to $736$, $1000$, $1000$, $800$, $1440$, $1056$, $1056$ for CTW1500, Total-Text, Inverse-Text, MSRA-TD500, and ICDAR 2015, ReCTS, and VinText, respectively. Following [liu2021abcnet, ye2023dptext, wang2022tpsnet], we adopt IoU@0.5 as the evaluation metric for text detection across all datasets. The evaluation metric for text spotting is consistent with [liu2020abcnet, huang2022swintextspotter, huang2023estextspotter]. The model is trained distributively on 4 NVIDIA RTX 3090 GPUs with a batch size of 4 per GPU, and all the inference speeds listed in the tables are tested on a single NVIDIA RTX 3090 GPU. When using the $250k$ iteration pre-training schedule for text spotting, the whole training process takes about 2.5 days.
@@ 4.3.1 Different Data Representations
TABLE III: Robustness comparison of FMS and SVD on CTW1500 under different ratios of artificial label noise.
Noise Method Detection E2E IoU
@@
5% SVD 84.9 88.1 86.5 67.1 $95.0^{(\downarrow 3.0)}$
FMS 85.5 88.5 87.0 67.7 $\mathbf{97.6}^{(\downarrow 0.5)}$
10% SVD 84.8 87.8 86.3 67.0 $94.4^{(\downarrow 3.6)}$
FMS 85.8 87.9 86.8 67.5 $\mathbf{97.4}^{(\downarrow 0.7)}$
20% SVD 84.5 87.5 86.0 66.7 $93.6^{(\downarrow 4.4)}$
@@ 4.3.3 Different Low-Rank Subspace Projections
FMS is theoretically more robust than traditional SVD to the outliers often found in real-world annotations. We conduct two experiments to verify this.
First, to validate the necessity of a robust method against the annotation errors inevitably introduced in real-world datasets, we create corrupted versions of the CTW1500 training set by injecting varying levels of spike noise (5%, 10%, and 20%), where one to five random vertices are drastically displaced. We then learn the LRA basis from this noisy data. As shown in Table III, the $\ell_{2}$-based SVD is highly sensitive to outliers; even at a low noise level of 5%, its representation quality (IoU) drops significantly by 3.0%. As the noise ratio increases to 20%, SVD suffers an additional 1.4% degradation in IoU. In contrast, our FMS method exhibits better resistance to increasing noise levels. It maintains a high IoU (97.2%) and a stable detection F-measure (86.8%) even under severe corruption (20% noise). This confirms that FMS extracts the intrinsic shape structure with less distortion caused by outliers, suggesting its potential for robust learning from noisy pseudo-labels in semi-supervised settings.
Second, we compare their performance on geometrically complex shapes using the highly curved text subset from [zhu2021fourier]. As shown in Table IV, FMS builds a higher quality projection. Crucially, at a low dimension of $M=6$, FMS achieves a significant IoU gain of 2.5% over SVD. This indicates that the principal components derived by FMS capture the essential geometric primitives of curved text more compactly than those from SVD. This representation advantage translates to performance gains of 0.7% and 0.8% in detection and spotting F-measures, respectively. When the dimension increases to 14, FMS continues to outperform SVD by 0.5% in IoU, with increases of 0.4% and 0.3% in F-measure for text detection and spotting, respectively. However, when the dimension increases to 22, although both FMS and SVD achieve excellent representation quality, consistent with Table II, there is no improvement in detection performance, as the difficulty of regressing high-frequency components outweighs the marginal geometric gains.
Overall, we adopt FMS as the low-rank subspace projection method due to its superior robustness against outliers and higher representation quality with compact dimensions compared to SVD.